A projectile is fired in a way that its horizontal range is equal to three times its maximum height .What is the angle of projection?


1
Expert's answer
2022-05-03T14:21:14-0400
"R=3H"

"\frac{v_0^2\sin2\theta}{g}=3\frac{v_0^2\sin^2\theta}{2g}"

"\tan\theta=\frac{2}{3}"

"\theta=33.4^\circ"